OpenStudy (anonymous):

"A Spherical tank of radius 10 is full of liquid of density (omega) lb/ft^3. Find the work W done by pumping the liquid to the top of the tank." Here's what I know: W=F*d. radius is 10 feet, and since the tank is half full, I can probably assume distance d would also be 10 feet. I also know the volume of the cylinder, which is pi*r^2*h. Using integration via cylinders, I's probably assume h is either dx or x. If it's dx (thickness), then the displacement d would actually be x, at which point I'm integrating from 0 to 10 (setting center of sphere as the origin).

OpenStudy (anonymous):

This would mean that my answer would be something along the lines of... \[\int\limits_{0}^{10} \pi(10)^2xdx\] My question is, did I get anything wrong here?

OpenStudy (anonymous):

Oh wait, forgot the omega: \[\int\limits_{0}^{10} \pi(10)^2\omega*xdx\]

OpenStudy (anonymous):

SO yeah, did I get any part of that wrong?

OpenStudy (anonymous):

I'm studying for an exam, so my grade could potentially be riding on whether I get confirmation on this.